OpenStudy (anonymous):

Given f(x)=(3x-1)^2 Find the equation of the line tangent to the graph x=2, and find the value of x where the tangent line is horizontal

OpenStudy (anonymous):

http://bit.ly/QU4mW2

jimthompson5910 (jim_thompson5910):

f(x) = (3x-1)^2 f ' (x) = 2(3x-1)^1 ... use the power rule now you need to multiply that with the derivative of the inside, so you need to derive 3x-1 to get 3, then you multiply 3 by that last expression on the right side to get f ' (x) = 3*2(3x - 1)^1 which simplifies to f ' (x) = 6(3x - 1) f ' (x) = 18x - 6

jimthompson5910 (jim_thompson5910):

now plug in x = 2 into f ' (x) and tell me what you get

OpenStudy (anonymous):

30

jimthompson5910 (jim_thompson5910):

that's the slope of the tangent line at x = 2

jimthompson5910 (jim_thompson5910):

so m = 2

jimthompson5910 (jim_thompson5910):

I meant, m = 30

jimthompson5910 (jim_thompson5910):

now we need to find the y coordinate of the point on f(x) when x = 2 f(x) = (3x-1)^2 f(2) = (3(2)-1)^2 ... plug in x = 2 f(2) = (6-1)^2 f(2) = 5^2 f(2) =25

jimthompson5910 (jim_thompson5910):

so the point (2,25) is on the function f(x)

jimthompson5910 (jim_thompson5910):

we have enough to find the equation of the tangent line at x = 2 this is because we have a slope (m = 30) and a point the line goes through (2,25)

jimthompson5910 (jim_thompson5910):

y = mx + b y = 30x + b 25 = 30(2) + b solve for b

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

-35

jimthompson5910 (jim_thompson5910):

so your tangent line is y = 30x - 35

jimthompson5910 (jim_thompson5910):

tangent line at x = 2

jimthompson5910 (jim_thompson5910):

that answers the first part of the question

jimthompson5910 (jim_thompson5910):

f ' (x) = 18x - 6 represents the slope of the tangent line at ANY point on f(x)

jimthompson5910 (jim_thompson5910):

if the tangent line is horizontal, then the slope is 0

jimthompson5910 (jim_thompson5910):

so when we want to find what values of x make horizontal tangent lines, we just plug in f ' (x) = 0 and solve for x f ' (x) = 18x - 6 0 = 18x - 6 ... ... x = ??

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

.3

jimthompson5910 (jim_thompson5910):

but it's a bit off

OpenStudy (anonymous):

repeating

jimthompson5910 (jim_thompson5910):

good or x = 1/3

jimthompson5910 (jim_thompson5910):

at x = 1/3, the tangent line is horizontal

jimthompson5910 (jim_thompson5910):

finding horizontal tangents is useful for maximizing or minimizing a function so you can apply it to maximizing profit or minimizing costs

OpenStudy (anonymous):

well...that seems very easy

jimthompson5910 (jim_thompson5910):

yeah it's not too bad once you get to know it

OpenStudy (anonymous):

that seemed like algebra easy...like I was WAY overthinking it!

jimthompson5910 (jim_thompson5910):

well you're definitely using calculus ideas and rules, but in a way, yes it's like algebra